OpenStudy (anonymous):

what is the ratio of x to y ? 3/5x/x+2y A.-3 B.-1/3 C.1/3 D. 3

OpenStudy (anonymous):

@Captain_Page_Turner

OpenStudy (anonymous):

@JFraser @Elsa213

OpenStudy (anonymous):

Help please.

OpenStudy (anonymous):

@BlossomCake

OpenStudy (anonymous):

@coolman500

OpenStudy (anonymous):

what do you think

OpenStudy (anonymous):

um c.

OpenStudy (anonymous):

no it is D 3

OpenStudy (anonymous):

D.3

OpenStudy (anonymous):

oh can you explain how to get that answer i need to know how to do it too lol

OpenStudy (anonymous):

okay lol I useally do

OpenStudy (anonymous):

Oh okay

OpenStudy (anonymous):

3/5=x/(x+2y) multiply both side by x+2y x=3/5 x (x+2y) x=3x/5 + 6y/5 x=(3x+6y)/5 multiply both side by 5 5x=3x+6y subtract both side by 3x 2x=6y simplify both side by 2 x=3y divide both side by y x/y=3

OpenStudy (anonymous):

Okay thank you so much :)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

i gave a medal and became a fan btw